OpenStudy (anonymous):

An equation is shown below. 5x - y = 8 Part A: Explain how you will show all of the solutions that satisfy this equation. Part B: Determine three different solutions for this equation. Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

@phi @ganeshie8 @undeadknight26 @tHe_FiZiCx99

undeadknight26 (undeadknight26):

@arabpride @ganeshie8 @RadEn @radar @TuringTest @thomaster

OpenStudy (the_fizicx99):

K, um for part A its impossible to show all the solutions as the line goes on infinitely. For part B, plug in any random value for x and solve. Suppose I have, 9x - y = -2 ---------------- x = 2 9(2) - y = -2 18 -y = -2 -18 -18 -y = -20 /-1 /-1 y = 20 For part C, write the equation in y = mx + b and make an inconsistent system. For example. 9x - y = -2 ----> y = 9x + 2 an inconsistent system would be y = 9x -3

OpenStudy (anonymous):

Thank you so much :D your my hero right now i was so stumped!

OpenStudy (the_fizicx99):

:*